# SemTalk: Holistic Co-speech Motion Generation with Frame-level Semantic Emphasis

### v1

@@ 3.4 Generating Semantic-aware Sparse Motion
The Sparse Motion Generation (Fig. 3 b) in SemTalk adds expressive, context-specific gestures to the rhythm-aligned base motion by incorporating sparse but meaningful semantic cues drawn from speech content and emotional tone. By separating rhythm and semantics, this stage enhances gesture generation with contextually nuanced movements, focusing on moments of high semantic relevance.
Semantic Speech Encoding. To capture the breadth of semantic cues in speech, Semantic Emphasis Learning combines three types of information: frame-level text embeddings $\phi_{l}$, sentence-level features $\phi_{g}$ from the CLIP model [38], and emotion features $\phi_{e}$ from the emotion2vec model [34]. Together, these features form a comprehensive semantic representation $f_{t}$ that reflects both the content and emotional undertones of speech, enabling SemTalk to activate motions that are sensitive to nuanced semantic cues.
Semantic Emphasis Learning. Semantic Emphasis Learning in SemTalk is designed to capture sparse, semantically rich gestures that enhance motion by focusing on frames of high contextual relevance. Learning semantic-aware motion separately from rhythm-based gestures enables SemTalk to highlight meaningful frames that enrich gesture expressiveness without disrupting the underlying base motion.
The process begins by generating fused semantic features $f_{t}$, combining local and global cues from text, speech, and emotion embeddings to help identify frames needing semantic emphasis. Using $f_{t}$, SemTalk computes a semantic score $\psi$, which highlights frames with high semantic value, while reducing attention to rhythm-only frames. $\psi$, informed by broader context, filters back semantic features $f_{t}$ to activate frames with significant relevance, ensuring that the model emphasizes frames aligned with specific communicative intentions. Supervision of $\psi$ through a classification loss $\mathcal{L}_{cls}^{G}$ based on semantic labels further refines the model’s ability to identify key frames. To propagate semantic emphasis cohesively across different body regions, we use Coarse2Fine Cross-Attention Module similar to Sec. 3.3 to get sparse semantic motion $f_{s}$. In this stage, the emphasis in Sparse Motion Generation is placed on body motion, as these are more critical for conveying semantic meaning in co-speech interactions.
@@
Figure 4: Comparison on BEAT2 [30] Dataset. SemTalk* refers to the model trained solely on the Base Motion Generation stage, capturing basic rhythmic alignment but lacking detailed semantic gestures. In contrast, SemTalk successfully emphasized sparse yet vivid motions. For example, when the phrase “my opinion” is spoken, SemTalk-driven characters raise both hands and then make the gesture of extending their index finger to emphasize the statement. When speaking “never tell,” our model produces a clear, repeated gesture to match the rhythm of each word, capturing the intended emphasis.
@@ 3.5 Semantic Score-based motion fusion
The Adaptive Fusion stage (Fig. 3 c) in SemTalk seamlessly integrates sparse semantic-aware motion into the rhythmic-related base motion. By strategically enhancing frames based on their semantic importance, it maintains a smooth and natural motion flow across sequences. For each frame $i$, the semantic score $\psi_{i}$ computed during the Sparse Motion Generation stage is compared to a threshold $\beta$. If $\psi_{i}>\beta$, the base motion’s latent code $q_{i}^{r}$ is replaced with the sparse semantic code $q_{i}^{s}$, effectively highlighting expressive gestures where they are most relevant; otherwise, $q_{i}=q_{i}^{r}$.
This selective replacement emphasizes semantically critical gestures while preserving the natural rhythmic base motion. By blending $q^{r}$ and $q^{s}$ based on semantic scores, SemTalk adapts to the expressive needs of the speech context while ensuring coherence. The result is a realistic motion sequence where important gestures stand out without disrupting the foundational rhythm. To ensure smooth transitions and consistency across all body parts, this frame-by-frame replacement prevents abrupt shifts and effectively emphasizes the most expressive frames.
@@ 4.1 Experimental Setup
Datasets. For training and evaluation, we utilize two primary datasets: BEAT2 and SHOW. BEAT2, as introduced in EMAGE [30], is an extension of the original BEAT dataset [29], which includes 76 hours data featuring 30 speakers. BEAT2 standardizes the BEAT into a mesh representation with paired audio, text transcripts, and frame-level semantic labels. Following [30], we focus on the BEAT2-standard subset, allocating data for 85%/7.5%/7.5% train/val/test. SHOW [47] contains 26.9 hours of high-quality, in-the-wild talk show videos featuring 3D body meshes at 30fps. Due to the absence of frame-level semantic labels in the SHOW dataset, we adapt the Sem-gate module from SemTalk, pre-trained on the BEAT2 dataset, to generate these labels. Following [47], we select video sequences longer than 10 seconds and divide the dataset into 80%/10%/10% train/val/test split.
Implementation Details. Our model is trained on a single NVIDIA A100 GPU. For both datasets, training runs for 200 epochs with a batch size of 64. We use the RVQ-VAE architecture [41] as the auto-encoder, with ResNet blocks in both encoder and decoder, downscaling by a factor of 4. The residual quantization employs 6 layers with a code dimension and codebook size of 256, and a dropout rate of 0.2, we predict the remaining five codes through five transformer layers from the first layer’s output. In Base Motion Learning, the temperature $\tau$ is set to 0.1. In Sparse Motion Learning, $\beta$ is set to 0.5. The training uses the ADAM optimizer with a learning rate of 1e-4. Inspired by [30], we first train with a four-frame seed pose, then gradually increase the masked joints and frames ratio from 0 to 40% over the initial 120 epochs, fixing it afterward.
Figure 6: Facial Comparison on the BEAT2 [30] Dataset. Our approach synchronizes facial expressions closely with phonetic and semantic cues in speech, generating natural lip movements that enhance clarity and expressiveness, as seen in words like ”always” and ”because,” where each syllable transition is smoothly captured in alignment with the audio input.
Metrics. We evaluate the generated body gestures using the FGD metric [49] to assess the distributional alignment between generated and GT, reflecting gesture realism. DIV [25] is measured by the average L1 distance across multiple gesture clips, indicating the range of gesture variation. To assess speech-motion synchrony, we employ BC [26], which measures the temporal alignment between gesture rhythm and audio beats. For facial expressions, we utilize two reconstruction metrics: vertex MSE [46] to quantify positional differences, and vertex L1 difference LVD [47] to gauge discrepancies between GT and generated facial vertices. Given the lack of explicit semantic consistency labels, we further conduct a comprehensive user study. We excluded ProbTalk [32] from our comparison because we were unable to obtain correct results during retraining using the publicly available code.
@@ 4.2 Qualitative Results
Qualitative Comparisons. We strongly encourage readers to view our demo video for a clear understanding of SemTalk’s qualitative performance. Overall, our method demonstrates superior alignment between speech and motion, generating more realistic, diverse, and semantically consistent motion than the baselines on both datasets. On the BEAT2 dataset (Fig. 4), TalkSHOW, EMAGE, and DiffSHEG all exhibit slight jitter effects. EMAGE displays noticeable jitters mainly in the legs and shoulders, while TalkSHOW suffers from instability throughout the body. DiffSHEG, designed primarily to focus on the upper body without considering the lower body holistically, generates slow motions with visible inconsistencies, especially at the boundaries of speech clips. Regarding motion diversity, DiffSHEG generates a broader range of gestures than EMAGE and TalkSHOW; however, EMAGE maintains a higher degree of naturalness. SemTalk surpasses all baselines by achieving the highest levels of both diversity and realism. Compared to the SemTalk* baseline, SemTalk generates gestures that are contextually appropriate, lively, and expressive. Although SemTalk* ensures rhythmic consistency, it lacks strong semantic expressiveness. SemTalk, by contrast, emphasizes key phrases, such as raising both hands to highlight ”dream job” or using a definitive pointing gesture for ”that is why,” thereby reinforcing the semantics. SemTalk* lacks expressiveness, but can still generate a well-formed, general base motion that is rhythmically consistent. With the semantic emphasis on frame-level motion, our SemTalk finally generates motions that are both rhythm-related and semantic-aware. This highlights the effectiveness of our rhythmic consistency learning and semantic emphasis learning.
@@
Figure 7: Qualitative study on semantic score.
On the SHOW dataset (Fig. 5), SemTalk shows more agile gestures than TalkSHOW, EMAGE, and DiffSHEG, when applied to unseen data. Our method captures natural and contextually rich gestures, particularly in moments of emphasis such as “I like to do” and “relaxing,” where our model produces lively hand and body movements that align with the speech content. Unlike baselines, which tend to generate repetitive or overly simplistic motions, SemTalk maintains a high degree of diversity and subtlety, producing gestures that vary naturally across different phrases.
Semantic Score. Fig. 7 illustrates the relationship between semantic emphasis and gesture intensity, represented by variations in the semantic score. To model this correlation smoothly, we extracted actual semantic scores from selected key frames, capturing trends in gesture emphasis. Words with high semantic significance, such as ”comes,” ”fantastic,” and ”captured,” align with peaks in the semantic score, highlighting moments where the model emphasizes expressive gestures. These peaks indicate points in the audio that prompt the generation of more dynamic and engaging body movements, creating a closer alignment between gesture intensity and the expected meaning of the speaking.
User study. We conducted a user study using 10 video samples, recruiting 25 participants with varied backgrounds to assess several aspects of generated motion, including realism, semantic consistency, motion-speech synchronism, and diversity. The study required participants to rank shuffled videos across different methods on these criteria. As shown in Fig. 8, our approach received dominant preferences across all metrics, particularly excelling in semantic consistency and motion diversity. This indicates that our model produces highly synchronized and contextually appropriate gestures that align well with the spoken content, outperforming baselines in delivering realistic and expressive motion that users find engaging and natural.
@@
[…]
- - - - 6.234 7.628 11.44 8.239 7.831
- - - $\surd$ 5.484 7.641 11.84 13.882 15.42
$\surd$ - - $\surd$ 4.867 7.701 12.38 6.201 6.928
$\surd$ $\surd$ - $\surd$ 4.526 7.751 12.83 6.215 6.997
- - $\surd$ $\surd$ 4.897 7.702 12.42 13.416 15.72
$\surd$ $\surd$ - - 5.831 7.758 11.97 6.587 7.106
$\surd$ - $\surd$ $\surd$ 4.397 7.776 12.49 6.100 6.898
$\surd$ $\surd$ $\surd$ $\surd$ 4.278 7.770 12.91 6.153 6.938
@@ 4.3 Quantitative Results
Comparison with Baselines. As shown in Tab. 2, on the BEAT2 dataset, SemTalk achieves superior results in FGD, MSE, and LVD, showing improved distribution alignment and reduced motion errors compared to previous methods. To ensure a fair comparison, we add a lower-body VQ-VAE to TalkSHOW, DiffSHEG, and SemTalk, following [30]. Notably, our approach reduces FGD significantly, indicating robust distribution matching, particularly in generating expressive and context-aware gestures. While TalkSHOW and EMAGE achieve competitive diversity (DIV) scores, SemTalk maintains a balance between high semantic relevance and natural motion flow.
@@
Table 2: Quantitative comparison with SOTA methods. Our SemTalk method consistently outperforms state-of-the-art baselines across both the BEAT2 and SHOW datasets. Lower values are better for FMD, FGD, MSE, and LVD. Higher values are better for BC and DIV. We report FGD×$\times 10^{-}1$, BC×$\times 10^{-}1$, MSE×$\times 10^{-}8$ and LVD×$\times 10^{-}5$ for simplify.
Ablation Study. We assess the impact of each component of our model on BEAT2 and present the results in Tab. 1, which reveals several key insights (more ablation results please see supplementary material) :
  • Rhythmic Consistency Learning (RC) is essential for achieving temporal coherence in motion sequences. By capturing and preserving consistent rhythmic patterns, RC enhances the natural flow of generated motions. This approach not only boosts performance on key metrics like FGD, LVD, and BC but also reduces the MSE, contributing to smoother and more realistic base motion.
@@
  • Coarse2Fine Cross-Attention Module (C2F) effectively refines motion details, improving BC, FGD, and DIV. When combined with RVQ and RC, C2F achieves the best MSE and LVD, highlighting its role in enhancing gesture realism and diversity through fine-tuning from coarse to refined levels.

### v3

@@ 3.4 Generating Semantic-aware Sparse Motion
The Sparse Motion Generation (Fig. 3 b) in SemTalk adds semantic-aware sparse motion to base motion by incorporating semantic cues drawn from speech content and emotional tone. By separating rhythm and semantics, this stage enhances motion generation by emphasizing contextually meaningful motion at key semantic moments.
Semantic Speech Encoding. To capture semantic cues in speech, Semantic Emphasis Learning combines three types of information: frame-level text embeddings $\phi_{l}$, sentence-level features $\phi_{g}$ from the CLIP model [39], and emotion features $\phi_{e}$ from the emotion2vec model [35]. These features form a comprehensive semantic representation $f_{t}$, together with audio feature $\gamma_{h}$, that reflects both the content and emotional undertones of speech, enabling SemTalk to activate motions that are sensitive to nuanced semantic cues.
Semantic Emphasis Learning. The process begins by generating $f_{t}$, combining local and global cues from text, speech, emotion embeddings and HuBERT features $\gamma_{h}$. Then, the sem-gate leverages multi-modal inputs to generate a semantic score, identifying frames that require enhanced semantic emphasis. The sem-gate in SemTalk refines keyframe motion by applying two forms of weighting methods $\mathcal{W}$: feature weighting $\mathcal{W}_{f}$ and loss weighting $\mathcal{W}_{l}$. Using $f_{t}$ and $\gamma_{h}$, SemTalk computes a semantic score $\psi$, which dynamically scales feature weighting—filtering back semantic features $f_{t}$ to activate frames with significant relevance, ensuring that the model emphasizes frames aligned with specific communicative intentions. Second, the loss weighting is applied by supervising $\psi$, with a classification loss $\mathcal{L}_{cls}^{G}$ based on semantic labels, further enhancing the model’s ability to identify key frames. The two weighting methods allow SemTalk to selectively enhance semantic gestures while suppressing uninformative motion, leading to more expressive co-speech motion.
@@
To ensure cohesive propagation of semantic emphasis across body regions, we employ the Coarse2Fine Cross-Attention Module, similar to Sec. 3.3. In this stage, we focuses solely on body motion, excluding facial movements, as body gestures play a more critical role in conveying semantic meaning in co-speech interactions.
@@
Figure 4: Concept comparison with LivelySpeaker [52]. (Top) LivelySpeaker generates semantic gestures with CLIP embeddings in SAG and refines rhythm-related gestures separately using diffusion, causing potential jitter. (Bottom) SemTalk integrates text and speech, uses a semantic gate for fine-grained control, and unifies rhythm and semantics for smoother, more coherent motions.
Discussion. Recently, LivelySpeaker [52] designs the Semantic-Aware Generator (SAG) and Rhythm-Aware Generator (RAG) for co-speech gesture generation, combining them through beat empowerment. While effective, key differences exist between LivelySpeaker and SemTalk, see Fig. 4. First, SAG generates gestures from text using CLIP embeddings, but bridging words and expressive gestures is challenging, causing jitter. SemTalk incorporates speech features (pitch, tone, emotion) alongside text and GT supervision for adaptive gestures. Second, LivelySpeaker applies global control, missing local semantic details, while SemTalk uses fine-grained, frame-level semantic control for subtle variations. Third, LivelySpeaker fuses SAG and RAG in separate latent spaces, leading to misalignment and inconsistencies. SemTalk jointly models rhythm and semantics in a unified framework, ensuring smoother transitions and coherence. We further compare SAG with our semantic gate in experiments.
Figure 5: Comparison on BEAT2 [31] Dataset. SemTalk* refers to the model trained solely on the Base Motion Generation stage, capturing rhythmic alignment but lacking semantic gestures. In contrast, SemTalk successfully emphasized sparse yet vivid motions. For instance, when saying “my opinion,” SemTalk generates a hand-raising gesture followed by an index finger extension for emphasis. Similarly, for “never tell,” our model produces a clear, repeated gesture matching the rhythm, reinforcing the intended emphasis.
@@ 3.5 Semantic Score-based motion fusion
The Adaptive Fusion stage (Fig. 3 c) in SemTalk seamlessly integrates semantic-aware sparse motion into the rhythmic-related base motion. By strategically enhancing frames based on their semantic importance, it maintains a smooth and natural motion flow across sequences. For each frame $i$, the semantic score $\psi_{i}$ computed during the Sparse Motion Generation stage is compared to a threshold $\beta$. If $\psi_{i}>\beta$, the base motion’s latent code $q_{i}^{r}$ is replaced with the sparse semantic code $q_{i}^{s}$, effectively highlighting expressive gestures where they are most relevant; otherwise, $q_{i}=q_{i}^{r}$.
This selective replacement emphasizes semantically critical gestures while preserving the natural rhythmic base motion. By blending $q^{b}$ and $q^{s}$ based on semantic scores, SemTalk adapts to the expressive needs of the speech context while ensuring coherence. Additionally, the convolution structure of the RVQ-VAE decoder ensures smooth transitions between frames, preserving motion continuity.
@@ 4.1 Experimental Setup
Datasets. For training and evaluation, we use two datasets: BEAT2 and SHOW. BEAT2, introduced in EMAGE [31], extends BEAT [30] with 76 hours of data from 30 speakers, standardized into a mesh representation with paired audio, text, and frame-level semantic labels. We follow [31] and use the BEAT2-standard subset with an 85%/7.5%/7.5% train/val/test split. SHOW [48] includes 26.9 hours of high-quality talk show videos with 3D body meshes at 30fps. Since it lacks frame-level semantic labels, we use the sem-gate from SemTalk, pre-trained on BEAT2, to generate them. Following [48], we select video clips longer than 10 seconds and split the data 80%/10%/10% for train/val/test.
Implementation Details. Our model is trained on a single NVIDIA A100 GPU for 200 epochs with a batch size of 64. We use RVQ-VAE [42], downscaling by 4. The residual quantization has 6 layers, a codebook size of 256 and a dropout rate of 0.2. We use five transformer layers to predict the last five layer codes. In Base Motion Learning, $\tau$ = 0.1; in Sparse Motion Learning, $\beta$ = 0.5 empirically. The training uses ADAM with a 1e-4 learning rate. Following [31], we start with a 4-frame seed pose, gradually increasing masked frames from 0 to 40% over 120 epochs.
Metrics.We evaluate generated body gestures using FGD [50] to measure distributional alignment with GT, reflecting realism. DIV [26] quantifies gesture variation via the average L1 distance across clips. BC [27] assesses speech-motion synchrony. For facial expressions, we use MSE [47] to quantify positional differences and LVD [48] to measure discrepancies between GT and generated facial vertices.
@@ 4.2 Qualitative Results
Qualitative Comparisons. We encourage readers to watch our demo video for a clearer understanding of SemTalk’s qualitative performance. Our method achieves superior speech-motion alignment, generating more realistic, diverse, and semantically consistent gestures than the baselines. As shown in Fig. 5, LivelySpeaker, TalkSHOW, EMAGE, and DiffSHEG exhibit jitter—EMAGE mainly in the legs and shoulders, while TalkSHOW affects the entire body. LivelySpeaker and DiffSHEG, which focus primarily on the upper body, produce slow and inconsistent motions, especially at speech clip boundaries. DiffSHEG improves gesture diversity over EMAGE and TalkSHOW, though EMAGE maintains greater naturalness. SemTalk surpasses all baselines in both realism and diversity. Compared to SemTalk*, SemTalk generates more expressive gestures, emphasizing key phrases (e.g., raising hands for “dream job” or pointing for “that is why”). While SemTalk* ensures rhythmic consistency, it lacks semantic expressiveness. By integrating frame-level semantic emphasis, SemTalk aligns motion with both rhythm and semantics, demonstrating the effectiveness of rhythmic consistency learning and semantic emphasis learning.
Figure 7: Facial Comparison on the BEAT2 [31] Dataset.
@@
Figure 8: Qualitative study on semantic score. Semantic score aligns with keywords, influencing gesture intensity.
On the SHOW dataset (Fig. 6), SemTalk shows more agile gestures than all baselines, when applied to unseen data. Our method captures natural and contextually rich gestures, particularly in moments of emphasis such as “I like to do” and “relaxing,” where our model produces lively hand and body movements that align with the speech content.
Semantic Score. Fig. 8 shows how semantic emphasis influences gesture intensity, with peaks in the semantic score aligning with keywords like ”comes,” ”fantastic,” and ”captured.” By extracting semantic scores from key frames, we track gesture emphasis trends. Furthermore, as shown in Fig. 9, SemTalk adapts to different emotional tones even when the text remains unchanged. This adaptability prevents overfitting to the text itself, allowing the model to generate gestures that vary according to the emotional delivery of the speech. The learned semantic score provides fine-grained, frame-level control, keeping gestures both rhythmically synchronized and semantically aligned in real time.
User Study. We conducted a user study with 10 video samples and 25 participants from diverse backgrounds, evaluating realism, semantic consistency, motion-speech synchrony, and diversity. Participants were required to rank shuffled videos across different methods. As shown in Fig. 10, our approach received dominant preferences across all metrics, especially in semantic consistency and realism.
Figure 9: Same words with different speech from the internet. “emo” represents different emotional tones extracted from speech. SemTalk can generate different motions, even when the text script is the same, preventing overfitting to the text itself.
@@ 4.3 Quantitative Results
Comparison with Baselines. As shown in Tab. 1, SemTalk outperforms previous methods on BEAT2, achieving lower FGD, MSE, and LVD, indicating better distribution alignment and reduced motion errors. For fairness, we follow [31] and add a lower-body VQ-VAE to TalkSHOW, DiffSHEG, and SemTalk. Notably, SemTalk significantly reduces FGD, ensuring strong distribution matching. While TalkSHOW and EMAGE achieve competitive diversity (DIV) scores, SemTalk balances high semantic relevance with natural motion flow.
@@
Table 1: Quantitative comparison with SOTA. SemTalk consistently outperforms baselines across both the BEAT2 and SHOW datasets. Lower values are better for FMD, FGD, MSE, and LVD. Higher values are better for BC and DIV. We report FGD×$\times 10^{-}1$, BC×$\times 10^{-}1$, MSE×$\times 10^{-}8$ and LVD×$\times 10^{-}5$ for simplify.
Sem-gate.  Tab. 2 highlights the effectiveness of sem-gate. Without sem-gate, the model fails to emphasize key moments. Randomized semantic scores led to poor performance by preventing meaningful frame distinction. Introducing a learned sem-gate even (w/o $\mathcal{W}$) significantly improves semantic alignment and classification accuracy. Refinement is further enhanced through weighting strategies: feature weighting $\mathcal{W}_{f}$ enhances motion emphasis, while loss weighting $\mathcal{W}_{l}$ improves FGD and overall accuracy. These results suggest that weighting methods enhance the accuracy of the semantic score and help the model prioritize important frames. The best results come from applying two weighting methods together, where frames with stronger semantic signals receive higher emphasis. We also compare sem-gate with LivelySpeaker’s SAG [52]. We find that replacing the Sparse Motion stage with SAG and substituting motion using GT semantic labels led to poor performance. SAG relies only on text-motion alignment, ignoring emotional tone, making it more prone to overfitting the text. In contrast, our sem-gate applies GT supervision with two weighting methods, achieving more accurate and stable semantic motion.
Ablation Study on Components. We assess the impact of each component of our model on BEAT2 and present the results in Tab. 3, which reveals several key insights (more ablation results please see supplementary material) :
  • Rhythmic Consistency Learning (RC) not only boosts performance on key metrics like FGD, LVD, and BC but also reduces the MSE, contributing to smoother and more realistic base motion.
@@
[…]
    w/o Sem-gate 4.893 7.702 12.42 -
    SAG (LivelySpeaker [52]) 4.618 7.682 12.45 -
    Sem-gate (Random $\psi$) 4.634 7.700 12.44 50.07
    Sem-gate (w/o $\mathcal{W}$) 4.495 7.633 12.26 72.32
    Sem-gate (w/ $\mathcal{W}_{f}$) 4.408 7.679 12.28 78.52
    Sem-gate (w/ $\mathcal{W}_{l}$) 4.366 7.772 11.94 77.83
    Sem-gate (ours) 4.278 7.770 12.91 82.76
    Table 2: Ablation study on Sem-gate. “Acc” denotes semantic classification performance on BEAT2. “w/o Sem-gate” means directly input $f_{t}$ and $\gamma_{h}$ without Sem-gate. “SAG (LivelySpeaker [52])” replaces the Sparse Motion Generation stage with LivelySpeaker’s SAG method. “Random $\psi$” assigns frame-level scores randomly. “w/o $\mathcal{W}$” applies the semantic gate but excludes frame-level weighting. “w/ $\mathcal{W}_{f}$” applies feature weighting. “w/ $\mathcal{W}_{l}$” applies loss weighting. (as mentioned in Sec. 3.4). Sem-gate (ours) integrates both the semantic gate and frame-level weighting to enhance emphasis.
[…]
    - - - - 6.234 7.628 11.44 8.239 7.831
    - - - $\surd$ 5.484 7.641 11.84 13.882 15.42
    $\surd$ - - $\surd$ 4.867 7.701 12.38 6.201 6.928
    $\surd$ $\surd$ - $\surd$ 4.526 7.751 12.83 6.215 6.997
    - - $\surd$ $\surd$ 4.897 7.702 12.42 13.416 15.72
    $\surd$ $\surd$ - - 5.831 7.758 11.97 6.587 7.106
    $\surd$ - $\surd$ $\surd$ 4.397 7.776 12.49 6.100 6.898
    $\surd$ $\surd$ $\surd$ $\surd$ 4.278 7.770 12.91 6.153 6.938
  • Coarse2Fine Cross-Attention Module (C2F) effectively refines motion details, improving BC, FGD, and DIV. When combined with RVQ and RC, C2F achieves the best MSE and LVD, highlighting its role in enhancing motion realism and diversity hierarchically.